OpenStudy (anonymous):

enter the quadratic equation in standard form x^2-12+4x=(x-2)^2+2

OpenStudy (p0sitr0n):

develop the square on the right side, then group all the values on the left side, you ll get Ax^2+Bx+C=0

OpenStudy (p0sitr0n):

iight so x^2 -12 +4x = x^2-4x+4+2 8x-12-6=0 8x-18=0 4x-9=0 which is your equation

OpenStudy (anonymous):

thanks